OpenStudy (anonymous):

To convert a length of 134 inches to yards, which ratio could you multiply by?

OpenStudy (whpalmer4):

you have 134 inches, but you want yards. Multiply inches by the following ratios: \[inches * \frac{1 {foot}}{12 {inches}} * \frac{1 {yard}}{3 {feet}}\] and after the units cancel out, you are left with yards. \[134 in*\frac{1 ft}{12 in}*\frac{1 yd}{3 ft} = 134 * \frac{1}{12}*\frac{1}{3} yd \approx 3.72 { }yd\] The numeric value of the ratio is 1:36. You have 1/36 as many yards as you do inches.